OpenStudy (anonymous):

Which of the following is the slope-intercept form of the equation -6x + y = 9 ? Select one: a. y = -6x-9 b. y = 6x+9 c. y = -6x+9 d. y = 6x-9

OpenStudy (bradely):

-6x + y = 9 Add 6x both sides 6x-6x + y=6x+ 9 y=6x+9 Option b iscorrect Source: http://www.mathskey.com/question2answer/ask

OpenStudy (anonymous):

b

OpenStudy (anonymous):

is not corret